I guess occupies a volume of 450 mL at 318 K. what will be the volume if the temperature is decreased to 268 K, if the pressure remains constant?


1
Expert's answer
2021-02-15T02:59:29-0500

At constant pressure,

T1/V1 = T2/V2

318/450 = 268/V2

V2 = 379 mL

Volume = [379] mL